# Nonlinear Concept Erasure: a Density Matching Approach

### v1

@@ E.1 GloVe
Figure 8: Gender erasure in GloVe with $\overline{\mathrm{L}}$ for different batch sizes used during training.
rank $a_{z}^{\mathrm{bin}}$downarrow (%) MDL $\uparrow$ (kBits) $a_{z}^{\mathrm{ter}}$downarrow (%) $A_{50\%}\uparrow$ WS-353 $\uparrow$
$\overline{\mathrm{L}}$
  ||||  $r=275$ full batch 88.8 $\pm$ 0.5 7.1 70.0 $\pm$ 0.4 0.88 0.72
@@
$\overline{\mathrm{L}}$ (cascaded)
  ||||  $r=275$ full batch 87.3 $\pm$ 0.4 8.0 68.1 $\pm$ 0.7 0.87 0.73
@@
Figure 9: Gender erasure in GloVe with $\overline{\mathrm{L}}$ vs. rank of the feature space post-erasure (left) and Erasure–Utility trade-off (right) for different batch sizes used during training. The vertical axis, erasure strength as evaluated by MDL, is shared between the right and left panels.
@@ E.2 Bias in Bios
Figure 10: Gender erasure in Bias in Bios with $\overline{\mathrm{L}}$ for different batch sizes used during training.
batch size $a_{z}$downarrow (%) MDL $a_{y}$uparrow (%)
$\overline{\mathrm{L}}$
  ||||  $r=200$ 8192 85.7 $\pm$ 0.4 142.3 76.8 $\pm$ 0.1
@@
$\overline{\mathrm{L}}$ (cascaded)
  ||||  $r=200$ 8192 57.3 $\pm$ 0.9 191.2 68.1 $\pm$ 0.6
@@
Figure 11: Gender erasure in Bias in Bios with $\overline{\mathrm{L}}$ vs. rank of the feature space post-erasure (left) and Erasure–Utility trade-off (right) for different batch sizes used during training. The vertical axis, erasure strength as evaluated by MDL, is shared between the right and left panels.
@@ E.3 DIAL
Figure 12: Race erasure in DIAL with $\overline{\mathrm{L}}$ for different batch sizes used during training.
batch size $a_{z}$downarrow MDL $\uparrow$ $a_{y}$uparrow DP $\downarrow$
$\overline{\mathrm{L}}$ (standard)
  ||||  $r=275$ 8192 86.5 $\pm$ 0.2 172.7 75.7 $\pm$ 0.1 0.19
@@
$\overline{\mathrm{L}}$ (cascaded)
  ||||  $r=275$ 8192 85.5 $\pm$ 0.2 185.3 74.8 $\pm$ 0.1 0.20
@@
Figure 13: Race erasure in DIAL with $\overline{\mathrm{L}}$ vs. rank of the feature space post-erasure (left) and Erasure–Utility trade-off (right) for different batch sizes used during training. The vertical axis, erasure strength as evaluated by MDL, is shared between the right and left panels.
@@ F.1 GloVe
Table 6: Gender erasure in GloVe with Cascaded $\overline{\mathrm{L}}$ for different number of kernels $M$ in the mixture.
rank $r$ $M$ $a_{z}^{\mathrm{bin}}$downarrow (%) MDL $\uparrow$ (kBits) $a_{z}^{\mathrm{ter}}$downarrow (%) $A_{50\%}\uparrow$ WS-353 $\uparrow$
@@ F.2 Bias in Bios
Table 7: Gender erasure in Bias in Bios with Cascaded $\overline{\mathrm{L}}$ for different number of kernels $M$ in the mixture.
rank $r$ $M$ $a_{z}$downarrow (%) MDL $a_{y}$uparrow (%)
@@ F.3 DIAL
Table 8: Race erasure in DIAL with Cascaded $\overline{\mathrm{L}}$ for different number of kernels $M$ in the mixture.
rank $r$ $M$ $a_{z}$downarrow MDL $\uparrow$ $a_{y}$uparrow DP $\downarrow$

### v2

@@ E.1 GloVe
Figure 8: Gender erasure in GloVe with $\overline{\mathrm{L}}$EOPARD for different batch sizes used during training.
rank $a_{z}^{\mathrm{bin}}$ $\downarrow$ (%) MDL $\uparrow$ (kBits) $a_{z}^{\mathrm{ter}}$ $\downarrow$ (%) $A_{50\%}\uparrow$ WS-353 $\uparrow$
$\overline{\mathrm{L}}$EOPARD
  $|$  $r=275$ full batch 88.8 $\pm$ 0.5 7.1 70.0 $\pm$ 0.4 0.88 0.72
@@
$\overline{\mathrm{L}}$EOPARD (cascaded)
  $|$  $r=275$ full batch 87.3 $\pm$ 0.4 8.0 68.1 $\pm$ 0.7 0.87 0.73
@@
Figure 9: Gender erasure in GloVe with $\overline{\mathrm{L}}$EOPARD vs. rank of the feature space post-erasure (left) and Erasure–Utility trade-off (right) for different batch sizes used during training. The vertical axis, erasure strength as evaluated by MDL, is shared between the right and left panels.
@@ E.2 Bias in Bios
Figure 10: Gender erasure in Bias in Bios with $\overline{\mathrm{L}}$EOPARD for different batch sizes used during training.
batch size $a_{z}$ $\downarrow$ (%) MDL $a_{y}$ $\uparrow$ (%)
$\overline{\mathrm{L}}$EOPARD
  $|$  $r=200$ 8192 85.7 $\pm$ 0.4 142.3 76.8 $\pm$ 0.1
@@
$\overline{\mathrm{L}}$EOPARD (cascaded)
  $|$  $r=200$ 8192 57.3 $\pm$ 0.9 191.2 68.1 $\pm$ 0.6
@@
Figure 11: Gender erasure in Bias in Bios with $\overline{\mathrm{L}}$EOPARD vs. rank of the feature space post-erasure (left) and Erasure–Utility trade-off (right) for different batch sizes used during training. The vertical axis, erasure strength as evaluated by MDL, is shared between the right and left panels.
@@ E.3 DIAL
Figure 12: Race erasure in DIAL with $\overline{\mathrm{L}}$EOPARD for different batch sizes used during training.
batch size $a_{z}$ $\downarrow$ MDL $\uparrow$ $a_{y}$ $\uparrow$ DP $\downarrow$
$\overline{\mathrm{L}}$EOPARD (standard)
  $|$  $r=275$ 8192 86.5 $\pm$ 0.2 172.7 75.7 $\pm$ 0.1 0.19
@@
$\overline{\mathrm{L}}$EOPARD (cascaded)
  $|$  $r=275$ 8192 85.5 $\pm$ 0.2 185.3 74.8 $\pm$ 0.1 0.20
@@
Figure 13: Race erasure in DIAL with $\overline{\mathrm{L}}$EOPARD vs. rank of the feature space post-erasure (left) and Erasure–Utility trade-off (right) for different batch sizes used during training. The vertical axis, erasure strength as evaluated by MDL, is shared between the right and left panels.
@@ F.1 GloVe
Table 6: Gender erasure in GloVe with Cascaded $\overline{\mathrm{L}}$EOPARD for different number of kernels $M$ in the mixture.
rank $r$ $M$ $a_{z}^{\mathrm{bin}}$ $\downarrow$ (%) MDL $\uparrow$ (kBits) $a_{z}^{\mathrm{ter}}$ $\downarrow$ (%) $A_{50\%}\uparrow$ WS-353 $\uparrow$
@@ F.2 Bias in Bios
Table 7: Gender erasure in Bias in Bios with Cascaded $\overline{\mathrm{L}}$EOPARD for different number of kernels $M$ in the mixture.
rank $r$ $M$ $a_{z}$ $\downarrow$ (%) MDL $a_{y}$ $\uparrow$ (%)
@@ F.3 DIAL
Table 8: Race erasure in DIAL with Cascaded $\overline{\mathrm{L}}$EOPARD for different number of kernels $M$ in the mixture.
rank $r$ $M$ $a_{z}$ $\downarrow$ MDL $\uparrow$ $a_{y}$ $\uparrow$ DP $\downarrow$